OpenStudy (anonymous):

differentiate y = ln(sin(lnx)) can you give me a hint of how to start? get rid of ln by raising the whole thing to the power of e?

OpenStudy (anonymous):

dy/dx=(dln(sin(lnx))/dsin(lnx))*(dsin(lnx)/d(lnx))*(dlnx/dx)) = cos(lnx)/xsin(lnx)

OpenStudy (anonymous):

there is no need to raising it to the power of e just use chain rule